# MapFormer: Self-Supervised Learning of Cognitive Maps with Input-Dependent Positional Embeddings

### v2

@@ 3.3 Path Integration Can Be More Easily Computed in the Lie-Algebra Space
The corresponding Lie algebra of $\mathbf{SO}(2)$, denoted $\mathfrak{so}(2)$, consists of all $2\times 2$ real matrices $A$ that are skew-symmetric ($A^{\top}=-A$). The Lie algebra $\mathfrak{so}(2)$ is a one-dimensional vector space, spanned by a single basis skew-symmetric matrix $S:=-S^{\top}$ (2D-rotations have a single degree of freedom). Thus, all elements of the Lie algebra can be expressed by a single scalar (the angular velocity) $\omega$ as $A=\omega S$. A finite rotation $R_{\theta}$ can then be seen as generated by integrating the infinitesimal angular velocities $\omega$ over a duration $\Delta_{t}$:
Intuitively, the skew-symmetric matrix $A=\omega S$ defines the instantaneous action direction $S$ and constant angular velocity $\omega$. The finite rotation $R_{\theta}\in\mathbf{SO}(2)$ is obtained by the exponential map, which performs a continuous integration of this velocity. Thus, the expression $R_{\theta}=\exp(\Delta_{t}A)$ shows that the final rotation yields a total angle of $\theta=\omega\Delta_{t}$, which defines the smooth path (geodesic) on the group manifold.
More generally, given the group of action-dependent matrices $W_{a}$’s (which, as we saw, needs to be a subgroup of the general linear group $GL(n,\mathbb{R})$), elements of the algebra, $A(t)$, can be expressed by a linear combination of generators $A_{i}$’s scaled by parameters $t_{i}$’s (the ‘integration times’, which can be seen as coordinates along the different directions defined by the generators). If surjective, all group elements $W_{a}$’s can then be expressed using the exponential map (fig.1):
For the special case where the generators $A_{i}$ commute (i.e., the Lie bracket is zero), as in the simple $\mathbf{SO}(2)$ example, the exponential of the sum equals the product of the exponentials: $W_{a}=\prod_{i=1}^{K}\exp(t_{i}A_{i})$. This shows again the benefit of the Lie algebra: it transforms the non-linear composition (multiplication) of the group into linear path integration (addition) within the algebra, thereby facilitating efficient learning and parallelization in the model.
3.4 MAmPa: Learning Cognitive Maps with block-diagonal Mamba Models
In this section, we propose a Lie-theoretic interpretation of State-Space Models (SSMs), which provides an interpretation of modern SSMs as models that implicitly perform path integration and can thus learn cognitive maps under certain conditions. This framework reveals a substantial limitation of current state-of-the-art SSMs, such as Mamba and Mamba2 gu2024mambalineartimesequencemodeling; dao2024transformersssmsgeneralizedmodels, to meet these conditions for learning cognitive maps. We thus propose a simple modification of the Mamba architecture, which we call MAmPa, for learning cognitive Maps with Mamba models.
Recently, structured state-space models (S4) gu2022efficientlymodelinglongsequences have gained popularity over Transformers as they allow parallelized training of sequences while keeping a linear inference cost. These models are inspired by continuous state space models, where a continuous signal $x(t)\in\mathbb{R}$ is mapped to $y(t)\in\mathbb{R}$ via a hidden variable $h\in\mathbb{R}^{n}$:
$h^{\prime}(t)$ $=Ah(t)+Bx(t)$ (7)
$y(t)$ $=Ch(t)$
$h_{t+1}$ $=\overline{A}h_{t}+\overline{B}x_{t+1}$ (8)
$y_{t+1}$ $=Ch_{t+1}$
where $\overline{A}=\exp(\Delta\cdot A)$ and $\overline{B}=(\Delta\cdot A)^{-1}\left(\exp(\Delta\cdot A)-I\right)\Delta B$. This discretization step is fundamental, as it explains how an exponential map can arise from the integration of infinitesimal recurrent transformations over a time interval $\Delta\in\mathbb{R}$.
$\overline{A}_{t}:=\exp(\Delta_{t}A)$ (9)
The exponential relationship between $A$ (the continuous domain matrix) and $\overline{A}$ (the discrete domain matrix) directly parallels the relationship between the element $A$ of the Lie algebra and the weight matrix $W_{a}$ in the corresponding Lie group. In the rotation example, since the Lie-algebra $\mathfrak{so}(2)$ of skew-symmetric matrices gives rise to the Lie-group of rotation matrices $\mathbf{SO}(2)$, $A:=\omega S$ is the infinitesimal generator and $W_{a}$ is the finite rotation over duration $\Delta_{t}$: $W_{a}=\exp(\Delta_{t}A)$. By identifying the discrete-time SSM matrix $\overline{A}$ with the action-dependent matrix $W_{a}$, we propose a new theoretical framework to analyze the capacity of Mamba to learn action-dependent matrices and cognitive maps. In this framework, Mamba can be interpreted as learning action-dependent matrices $\overline{A}_{t}$, which are dynamically created via an exponential map. As we saw, the exponential map integrates infinitesimal steps via $A$ (the direction and velocity in the Lie algebra space) over a varying time interval $\Delta_{t}$, giving a finite transformation $\overline{A}$ (the element of the Lie group). In other words, Mamba implicitly computes path integration of action $A$ over a duration defined by the learned parameter $\Delta_{t}$.
Given this theoretical framework, it is evident that current versions of Mamba are incapable of learning cognitive maps that require, for example, learning rotations. In Mamba, $A$ was chosen as diagonal, to reduce computational cost, but a diagonal action-dependent matrix cannot learn rotations. This is since $\overline{A}$ needs to be similar to a block-diagonal matrix, with blocks of size 2 (eq. 2 & 3). Mamba was shown to be able to learn gating problems, which increases the performance of Mamba on various structure-dependent tasks by ignoring irrelevant information, however, Mamba with a diagonal constraint is not expressive enough to represent spatial cognitive maps as Lie-group theory shows. In this work, we modify $A$ in Mamba to be block-diagonal and, in particular, to blocks of size 2 (for learning rotations) and show that this modification increases the capacity of the model to learn cognitive maps.
3.4.1 MAmPa as a Model of Episodic Memory (EM) and Working Memory (WM)
In Whittington2025, the authors introduce two related models of episodic (EM) and working memory (WM) that both learn cognitive maps using RNNs with input-dependent matrices. Here, we explain why selective, discrete State-Space Models (SSMs) can be viewed as EM and WM models. First, note that discrete SSMs can be regarded as RNNs with no activation function. Second, selective SSMs learn input-dependent matrices via an exponential map (eq. 8), meaning that their inner state captures the task’s latent structure, as in EM/WM models. Therefore, EM and WM models can naturally be regarded as selective SSMs.
SSM as a model of episodic memory
The EM-RNN model behaves exactly like TEM Whittington770495, factorizing structure and content in two distinct neural populations, but only updates its position with an input-dependent RNN $p_{t}:=W_{a_{t}}p_{t-1}$. It then queries a Hopfield memory module ramsauer2021hopfieldnetworksneed with its updated position to retrieve passed observations. The RNN in TEM can be seen as an SSM with subtle differences: (1) it uses no external input, i.e. $\forall t,x_{t}=0$; (2) a learnable embedding $p_{*}$ representing the cognitive map’s zero coordinate must be introduced to avoid the state being always null; (3) SSM’s inner state must be projected back to its final state ($y_{t}=Ch_{t}$, eq.8), which implies in this case that the final position $q_{t}:=Cp_{t}$ can be obtained with an optional learnable projection $C$; (4) as in TEM, the memory module is queried with the conjunction $g_{t}$ of what has been seen and where it was observed (fig. 1(c)):
$q_{t}$ $=Cp_{t}=CR_{\theta_{0\to t}}p_{*}$
$g_{t}$ $=\text{vec}(x_{t}^{\top}.q_{t})$
Setting $C:=\mathbf{1}_{n}$ implies that $q_{t}=p_{t}$, which is analogous to TEM/EM RNNs. Moreover, compared to RNN based models, this formulation allows to compute path integration directly in the Lie algebra ($A\sum_{0}^{t}\Delta_{s}$) and avoid sequential matrix multiplication.
SSM as a model of working memory
WM-RNNs directly updates the internal state with input-dependent matrices, which is precisely what selective SSMs do. As such, current selective SSMs with the appropriate matrix structure (skew-symmetric in the case of navigation, eq. 3) naturally model working memory (fig. 1(d)).
In the previous section, analyzing SSMs through Lie-group theory has allowed us to understand their current limitations, and that one only needs to switch to block-diagonal matrices to theoretically be able to learn a cognitive map. However, previous work decided to keep diagonal matrices because it dramatically reduces the computational load. By treating $A$ as a vector, one can avoid computing a matrix exponential and also dramatically reduce the memory footprint of a parallel scan, by computing vector-vector multiplications instead of matrix-vector’s. In Mamba2 dao2024transformersssmsgeneralizedmodels, it was further reduced to a scalar $A:=a\in\mathbb{R}$, which allows them to implement a fully parallelized version of selective SSMs but looses the ability to learn cognitive maps, as we have seen above. This highlights a fatal limitations of parallel-selective SSMs: they can be parallelized at the cost of loosing any hope of learning a cognitive map (see sec. 10.1 for further details). This is the reason why we decided to move away from SSMs and adapt Transformers to learn cognitive maps, as models of episodic (MapFormer-EM) and working memory (MapFormer-WM).
@@ 3.5 MapFormers: Path-Integrating Transformers for learning Cognitive Maps
In this section, we introduce two new Transformer-based architectures, which we call MapFormers, designed to learn path integration and, consequently, cognitive maps. Like RNNs and SSMs, the two proposed architectures primarily differ in how structural and content information are encoded (fig. 1c&d, fig. 2). The first architecture extracts and represents structural information separately from the content, using absolute positional embeddings to form a cognitive map; this is akin to a model of episodic memory Whittington770495. The second architecture, conversely, represents the sequential information of both structure and content jointly through relative positional encoding. This can be seen as a model of working memory that utilizes different neural-activity subspaces to store information in distinct memory slots Whittington2025. Crucially, the fundamental mechanism in both MapFormers is the use of input-dependent matrices to update positional encoding.
As we saw, a key aspect of cognitive maps that explains their contribution to OOD generalization is the factorization of content and structure in the input. In Transformers, this factorization can naturally arise when using input-dependent positional embeddings. To see this, we will start by reviewing the Transformer architecture.
The attention mechanism in Transformers is what has allowed the parallel processing of sequences, and therefore the ability to scale training on large scale datasets. An input sequence $X\in\mathbb{R}^{t\times d}$ is projected into keys, queries and values, $Q,K,V\in\mathbb{R}^{t\times d}$, and token mixing is achieved via a weighted sum of the values based on the key-query similarity matrix $A\in\mathbb{R}^{t\times t}$:
$\textbf{Att}(Q,K,V)=\mathrm{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V$ (11)
Since the attention mechanism is agnostic to the position in the sequence (and therefore invariant to permutation of the input), for many tasks, such as language modeling, positional information needs to be provided. There are two general approaches to introduce positional information – absolute positional and relative positional embedding. In the case of absolute positional embeddings, the embeddings of input token $x_{i}$’s is combined with positional information via: $x_{i}\mapsto x_{i}\odot p(i)$, where $\odot$ can be any binding operation, usually ++ or $\times$. These absolute positional embeddings $p_{i}$’s $\in\mathbb{R}^{d}$ can be either learnable embeddings or fixed (e.g., vectors of sines and cosines at different frequencies NIPS2017_3f5ee243). In the case of relative positional encoding, a function $e(i,j)$ is directly applied to the attention mechanism, to represent the relative position between a query $q_{i}$ and key $k_{j}$: $a_{ij}\mapsto a_{ij}\odot e(i,j)$.
@@
3.5.1 MapFormers-EM: MapFormers with Absolute Positional Embeddings as Models of Episodic Memory (EM)
We begin by introducing the first architecture of MapFormers, which makes use of absolute positional embeddings to learn a cognitive map, inspired by the links established between TEM and transformers (TEM-t whittington2022relatingtransformersmodelsneural). The key idea is to encode positions in a cognitive map using the keys and queries of the positional embeddings $p_{t}$’s, which are separated by design from the token embeddings $x_{t}$’s. For encoding a cognitive map, where the position in the map depends on previous actions in the input, fixed positional embeddings (e.g., vectors of sines and cosines at different frequencies NIPS2017_3f5ee243) would clearly not suffice since they can only represent a fixed structure (linear order), and therefore positional embeddings that are learned from the input are needed, i.e., input-dependent positional embeddings. This generalizes the notion of positional embeddings, which can now be understood as representations of the model’s position in an abstract, context-dependent cognitive map, and not simply position in the sequence.
To see how a cognitive map can be implemented via positional embeddings, we illustrate it with the example of navigation in 1D. We model the 1D movement on a circle rather than a line, since this enforces periodicity and prevents the positional representation from diverging (as opposed to exponential gating used in Mamba who’s inverse diverges towards infinity), allowing the model to naturally handle boundary conditions. The model’s objective is to track the absolute position $p_{t}$ at step $t$, which is updated after each action via a rotation matrix $R_{\theta}$. This position $p_{t}$ must be the result of path integration: the summation of all preceding actions in the input sequence. The key realization here is that path integration is simpler to compute in the Lie Algebra vector space, rather than in the Lie Group space (where positions are updated by costly matrix multiplication; section 3.3). In the linear Lie Algebra space, path integration is learned by simply adding up the small integration durations $\Delta_{t}$’s for each time step. Since this mechanism relies on a parallelizable summation operation, the Transformer architecture is ideally suited for this task, ensuring fast and efficient training and path computation.
In practice, path integration of the various $\Delta_{t}$’s across time is performed using a cumulative sum (cumsum) (fig.2(b)). For example, in the case of 1D navigation, each action token in the sequence is first projected onto a one-dimensional vector (a scalar in this case), which represents the integration time $\Delta_{t}$. The cumsum operation then computes, for each token, the estimated cumulative integration time $\sum_{i\leq t}\Delta_{i}$ (the total path length traveled so far). To compute the corresponding cumulative rotation angle ($\theta_{t}$) for each token, the cumulative integration time needs to be multiplied by the angular velocity $\omega$, a parameter learned by the model. In fact, to encode the full position in the cognitive map, the model learns a set of these frequencies ($\omega$’s), which jointly estimate the position in the map. These $\omega$’s are analogous to the different frequencies ($\omega=2\pi f$) used in standard fixed positional embeddings. Given the cumulative integration time and the learned frequencies, the rotation angle at time $t$ is computed. The next step is the re-entry into the Lie Group space: using the exponential map, we translate the linear path integration result from the Lie Algebra (addition) back into the action-dependent rotation matrices (elements of the Lie Group). This performs the required operation without needing to compute expensive matrix multiplication at every step by exploiting the property of the exponential map: $\exp\left(\sum_{i}\Delta_{t_{i}}A\right)=\prod_{i}\exp(\Delta_{t_{i}}A)$. In practice, for navigation problems, the matrix exponentiation does not need to be computed because the resulting form of the rotation matrix is already known (eq. 3). The resulting $\theta_{t}$’s from the path integration can thus be simply introduced into the rotation matrices, which then multiply the keys and queries of the positional embeddings.
Finally, the keys and queries of the positional embeddings, $K_{P}$ and $Q_{P}\in\mathbb{R}^{t\times d}$, are the vectors that encode the position in the cognitive map. Their starting point $p_{*}$, encoded by $k^{p}_{*},q^{p}_{*}\in\mathbb{R}^{d}$, represents the zero coordinate of the cognitive map and is learned as another set of parameters by the model. For each token in the sequence, these initial zero-coordinate vectors are updated by multiplying them with the action-dependent matrices $W_{a}$’s (the rotation matrices $R_{\theta}$’s in the case of navigation). These matrices represent the cumulative effect of the path integration, computed using the exponential map as described above. That is:
$p_{t}:=W_{t}p_{t-1}=\left(\prod_{s=1}^{t}W_{s}\right)p_{*}=\exp\left(\omega S\sum_{s=1}^{t}\Delta_{s}\right)p^{*}=R_{\theta_{0\to t}}p_{*}:=\mathbf{PI}_{0\to t}$ (12)
As we saw, the path integration thus avoids matrix multiplication and is replaced by the cumsum $\sum_{s}\Delta_{s}$, before computing matrix exponential, allowing scaling of this operation to much longer sequences by computing it in parallel. This contrasts with TEM-t, that sequentially updates its positional embeddings, loosing the parallel processing abilities of transformers.
The updated positional embeddings ($K_{P}$ and $Q_{P}$) constitute a ‘neural population’ encoding of the cognitive map. At each time step $t$, the resulting vectors $k^{p}_{t}$ and $q^{p}_{t}$ encode the path-integrated position $\mathbf{p}_{t}$. This positional module operates autonomously, equivalent to an SSM driven solely by internal dynamics (eq.10). The attention of MapEM mimics the role of the Hopfield module in EM-SSMs/RNNs, used to retrieve passed memories with the conjunction of what has been seen and where: $g_{t}=\text{vec}(x_{t}^{\top}.p_{t})$ (fig.2(b)). We extend this logic, such that the final positional keys and queries $k_{i}^{g}=\text{vec}({k^{x}_{i}}^{\top}.k_{i}^{p})$ and $q_{j}^{g}=\text{vec}({q^{x}_{j}}^{\top}.q_{j}^{p})$ reflect this conjunction.
Since $\forall i,j:\langle g_{i},g_{j}\rangle=\langle{q^{x}_{j}}^{\top}.q^{p}_{i},{k^{x}_{i}}^{\top}.k^{p}_{j}\rangle=\langle{q^{x}_{j}},k_{j}^{x}\rangle\langle q^{p}_{i},k^{p}_{j}\rangle$, we can avoid the outer products $Q_{X}^{\top}Q_{P}$ and $K_{X}^{\top}K_{P}$ by computing content and position attentions $A_{X}=\textbf{Att}(Q_{X},K_{X})$ and $A_{P}=\textbf{Att}(Q_{P},K_{P})$ separately (still in parallel by concatenating the heads), such that $A_{P}$ acts as an attention mask on $A_{X}$:
$\left(\textbf{Att}(Q,K)\odot\textbf{Att}(Q_{P},K_{P})\right)V=\left(\mathbf{A}_{X}\odot\mathbf{A}_{P}\right)V$ (13)
@@ 3.5.1 MapFormers-EM: MapFormers with Absolute Positional Embeddings as Models of Episodic Memory (EM)
3.5.2 MapFormers with Relative Positional Embeddings as a Model of Working Memory
As explained before, it was suggested that the brain encodes the structure of sequences, builds cognitive maps and stores them using two different systems: (1) Episodic Memory (EM), which stores a conjunction of structure and content information in long-term memory, as we saw previously, and (2) Working Memory (WM), which encodes sequential information without an explicit separation of the content from the learned map structure. Instead, WM employs dynamic encoding mechanisms to update structured activity slots.
In Transformers, position encoding can be achieved in a separate neural population, which mimics EM models with absolute positional embeddings, or by directly rotating the keys and queries to encode relative distances, like in WM models that directly update the state. The most popular positional encoding method is called RoPE (Rotary Positional Embeddings, su2023roformerenhancedtransformerrotary), where each key and query is rotated by $2\times 2$ block-diagonal rotation matrices $q_{i},k_{j}\mapsto R(i)q_{i},R(j)k_{j}$ such that the attention mechanism computes:
$a_{ij}:=\text{RoPE}(q_{i},k_{j}):=(R(j)k_{j})^{T}\cdot R(i)q_{i}=k_{j}^{T}R(i-j)q_{i}$ (14)
Since the transpose of a rotation matrix is also its inverse, the key-query dot product naturally relates to a rotation dependent on the distance iji-j. This property suggests a potential for implementing path integration (PI) in the form $\overline{q}_{j}^{\top}\cdot\overline{k}_{i}=q_{j}^{\top}R_{\theta_{PI_{i\to j}}}k_{i}$. However, the rotation matrices in standard RoPE are fixed, depending only on the positional distance iji-j. To learn the action-dependent matrices required for building a dynamic cognitive map, RoPE needs to be modified. This is achieved in MapFormers-WM by learning the rotation angle $\theta_{t}$ (and thus the matrix $R_{t}$) from the input context $x_{t}$, making it selective. Compared to simple gating, which only maintains or suppresses information, learned rotations allow tokens to interact in similar subspaces or ignore each other by rotating into orthogonal subspaces.
For efficient computation and fast parallelization, bot MapWM and MapEM implement linear path integration. This is achieved using cumulative summation (cumsum), as illustrated in fig. 2. During the training process, the model learns both the integration times ($\Delta_{t}$’s) and the angular velocities ($\theta_{t}$’s). In contrast to EM, the Working Memory (WM) approach results in the entanglement of structural and content information within the resulting positional embedding keys and queries ($K_{PI},Q_{PI}$), reflecting the integrated nature of sequential encoding in WM.
3.6 Relationship between MapFormers and MAmPa: Linear MapFormers are selective SSMs
In the previous section, we have introduced the exponential map in Transformers to allow them to build input-dependent matrices and learn cognitive maps. The Transformer formalism cannot naturally explain how an exponential map emerges like in SSMs, but in Mamba2 dao2024transformersssmsgeneralizedmodels, the authors show that linear transformers wang2020linformerselfattentionlinearcomplexity are actually SSMs. Starting from equation 8:
$h_{t}$ $=A_{t}h_{t-1}+B_{t}x_{t}$ (15)
after multiplying by $C_{t}$, we get the final equation:
$y_{t}=\sum_{s=0}^{t}C_{t}^{T}A^{\times}_{s\to t}B_{s}x_{s}$ (16)
$\Rightarrow$ $y=Mx;\quad m_{ij}=C_{j}^{T}A_{j}.A_{i+1}B_{i}$
If we choose matrix $A:=S$ to be a $2\times 2$ block-diagonal skew-symmetric matrix and $R_{t}:=\exp(\theta_{t}S)=\exp(\Delta_{t}\omega S)$ as in equation 5, we retrieve equation 14:
$m_{ij}$ $=q_{j}R^{\times}_{i\to j}k_{i}$ (17)
$=q^{T}_{j}R^{\times}_{0\to j}\left(R^{\times}_{0\to i}\right)^{-1}k_{i}$
$=\text{RoPE}(q_{j},k_{j})$
This proves that linear RoPE transformers are SSMs with a $\mathbf{2\times 2}$ skew-symmetric recurrent matrix, where $\theta_{t}$ controls the angle of rotation. In Mamba2, the authors define $A:=a\in\mathbb{R}$ as a scalar, giving $M_{ij}=\exp(-a\sum_{t=i+1}^{j}\Delta_{t})q_{j}^{\top}k_{i}$, which is analogous to the cumulative gating of CoPE (Contextualized Positional Encoding) golovneva2024contextualpositionencodinglearning, with a multiplicative and not additive binding operation. CoPE is an enhanced Transformer architecture with context-dependent positional encoding, that like Mamba compresses the relative key-query distances with a cumulative gating mechanism, and in both models, input selectivity allows them to focus or ignore certain tokens and solve structure-dependent tasks, but should not allow them to learn spatial cognitive maps.
This ends our unification of MAmPa and MapFormers with EM and WM RNNs. The formalism of SSMs allows us to understand how input-dependent matrices are created while computing path integration in parallel. Insights from Mamba2 allow us to unify relative positional embeddings with models of working memory, and complete the understanding of cognitive maps in transformers started in TEM-t whittington2022relatingtransformersmodelsneural. Finally, our analysis of SSMs through Lie-group theory demonstrates that diagonal matrices – necessary to reduce computation and scale SSMs – are not enough to learn spatial cognitive maps, leaving only MapFormers as realistic models of cognitive maps at scale.

### v3

@@ 3.2 Cognitive Maps and Additive Path-Integration
For example, in the case of 2D rotations, action-dependent matrices are rotation matrices $R_{\theta}$ that lie on the manifold of the special orthogonal group $\mathrm{SO}(2)$. Its Lie algebra $\mathfrak{so}(2)$ is one-dimensional, possessing a single generator matrix $S$ (the unit skew-symmetric matrix; $S^{T}=-S$), which spans the vector space of the algebra. Any matrix in $\mathfrak{so}(2)$ can be expressed as $A=\omega S$, where $\omega$ is the angular velocity. A finite rotation $R_{\theta}$ can then be generated by integrating the infinitesimal angular velocities $\omega$ over a duration $\Delta_{t}$. Mapformers can compute path integration at the level of the Lie algebra by summing the integration durations $\Delta_{t}$’s:
$W_{j\to i}:=\exp\left(\omega S\sum_{t=j}^{i-1}\Delta_{t}\right):=R(\theta_{j\to i}),$ (1)
Most of the computational cost thus boils down to a simple sum $\sum_{t=j}^{i-1}\Delta_{t}$. Computing the exponential map needs to be done only once, after the summation; furthermore, in the case of rotation, it is not strictly necessary, as the closed-form solution for the exponential map is already known to be the standard rotation matrix $R_{\theta}$ (sec. A.2,3,6&7).
In practice, in the case of navigation on a finite 2D grid, i.e. a torus, action-dependent matrices are $2\times 2$ block-diagonal rotations in $\mathrm{SO}(2)$, where each block is defined by a constant angular velocity $\omega_{b}$. They represent rotations at different scales, and are learned by the model during training, adapting themselves to the size of the grid – e.g., larger grids require a smaller $\omega$ to avoid wrapping around. Path integration can be computed along each of the $n_{b}$ blocks, in parallel.
Interestingly, the exponential map also appears following the discretization step of continuous State Space Models (SSMs) (Gu and Dao, 2024), by a step size $\Delta$. The continuous and discrete version of the SSM equations can thus be seen as describing dynamics occurring in a 1-dimensional Lie algebra and Lie group, respectively. Since Linear Transformers are equivalent to SSM (Dao and Gu, 2024), in this work, we mainly focus on $\mathrm{SO}(2)$, as it’s the only 1-dimensional compact Lie-group. Compactness is desired since a finite neural population can only represent a bounded cognitive map. This ensures commutativity, hence linear path-integration, and a closed-form solution of the exponential map (see sec. B.2 for non-abelian groups).
4 Models
In this section, we introduce two new Transformer-based architectures, which we call MapFormers, designed to learn path integration and, consequently, cognitive maps. The fundamental mechanism in both MapFormers is the use of input-dependent matrices to update positional encoding as in sec. 3.2, which, as explained, represents position in a context-dependent cognitive map instead of ordinal order used in standard Transformers.
Transformers achieve parallel and large-scale sequence processing via the attention mechanism. An input sequence $X\in\mathbb{R}^{l\times d}$ is projected into keys, queries and values, $Q,K,V\in\mathbb{R}^{l\times d}$, and token mixing is achieved via a weighted sum of the values based on the key-query similarity matrix $A\in\mathbb{R}^{l\times l}$:
$\textbf{Att}(Q,K,V)=\mathrm{softmax}\left(\frac{QK^{T}}{\sqrt{d}}\right)V$
Since the attention mechanism is permutation-invariant, positional information must be provided, either through absolute positional embeddings added to the input (Vaswani et al., 2017), or relative positional encoding to model the distance between keys and queries (Su et al., 2023).
The two proposed architectures (fig. 2(a) &  2(b)) follow this distinction: the first architecture extracts and represents structural information separately from the content, using absolute positional embeddings to form a cognitive map; this is akin to a model of episodic memory (Whittington et al., 2019). The second architecture, conversely, represents structure implicitly by rotating neural activity into distinct subspaces, akin to the models of working memory in Whittington et al. (2025).
To learn a cognitive map, the model has to identify that action tokens ($s_{t}=a_{t}$) update the internal position $p_{t}$, while ensuring that observations ($s_{t}=o_{t}$) leave the structure untouched. To optimize this process, the model learns the rotation duration rather than the matrix itself, as in eq. 1. Specifically, the model maps a token representation $x_{t}\in\mathbb{R}^{d}$ to a per-block integration duration $\{\Delta_{t_{b}}\}_{b}=f_{\Delta}(x_{t})\in\mathbb{R}^{n_{b}}$, and builds from it the final rotation angles $\theta_{t}=\{\omega_{b}\Delta_{t_{b}}\}_{b}$.
Then, rotations are aggregated in parallel along the temporal dimension $\theta_{0\to t}:=\mathrm{cumsum}_{t}(\theta_{t})$ to build the final block-diagonal, integrated rotation matrix $R_{\theta^{PI}}:=\{R_{\theta_{0}\to t}\}_{t\leq l}\in\mathbb{R}^{l\times d\times d}$ (fig. 2(a) &  2(b)). Here, $\omega\in\mathbb{R}^{n_{b}}$ represents the angular velocities on each rotation block, a parameter learned by the model during training (see sec. A.7 & A.8 for more details).
@@ 4.1 MapFormers: Path-Integrating Transformers for learning Cognitive Maps
MapEM: MapFormers with Absolute Positional Embeddings as Models of Episodic Memory (EM)
The first architecture of MapFormers makes use of absolute positional embeddings to learn a cognitive map (fig. 2(a)), inspired by the links established between TEM and transformers – TEM-t (Whittington et al., 2022). Once $R_{\theta^{PI}}$ is computed, one only needs to learn an embedding $p_{0}\in\mathbb{R}^{d}$, representing the origin of the cognitive map, such that:
$P:=R_{\theta^{PI}}P^{*};\quad P^{*}=[p_{0},\dots,p_{0}]_{l}\in\mathbb{R}^{l\times d}$
Attention is computed on $Q_{G}:=\{q^{g}_{t}=\mathrm{vec}(q_{t}^{\top}.p_{t})\}_{t\leq l}\in\mathbb{R}^{l\times d^{2}}$ and $K_{G}$, the conjunctions of what has been seen and where (fig.2(a)). This can be reframed as the element-wise product of content and position attentions $A_{X}=Q_{X}K_{X}^{T}/\sqrt{d}$ and $A_{P}=P.P^{T}/\sqrt{d}$, computed separately (still in parallel by concatenating the heads, see sec. A.7), such that $A_{P}$ acts as an attention mask on $A_{X}$:
$\textbf{Att}(Q_{g},K_{g},V)=\mathrm{softmax}\left(\mathbf{A}_{X}\odot\mathbf{A}_{P}\right)V$ (2)
This allows modeling Episodic Memory in a transformer that we call MapFormer-EM (or MapEM, in short, see sec. A.4 for details on EM models).
@@
MapWM: MapFormers with Relative Positional Embeddings as a Model of Working Memory
In (Whittington et al., 2025), Working Memory RNNs represent structure implicitly by shifting neural activity into structured activity slots with action-dependent matrices. This directly models the logic of RoPE (Su et al., 2023), where a fixed rotation matrix modeling relative distance is directly applied on the keys and queries, i.e. the state. In MapFormers-WM (MapWM; 2(b))), keys and queries are rotated by $R_{\theta^{PI}}$, which encodes their relative distance in the cognitive map. We further formalize the links between MapWM and models of working memory in sec. A.6.